# Also at ] The European Space Agency Centre for Earth Observation, Frascati, Italy

### v4

@@ II Methodology
At this point, We identify the characteristic scales in out model; the kernel size $\sigma$ and the active swim speed $v_{0}$, which yield the interaction time $\tau\equiv\sigma/v_{0}=3\text{ grid units}/0.5\text{ grid units/s}=6$ s, yielding the dimensionless units $\tilde{x}=x/\sigma$, $\tilde{t}=t/\tau$, and $\tilde{\omega}=\omega(\sigma/v_{0})$ (we subsequently drop the tilde). The system size becomes $\Lambda=2L/\sigma\approx 13$. A natural definition of the RFE is then afforded by $\tau$; we define the RFE as the center of mass of a particle’s trajectory during $\tau$, effectively course-graining the singularities:
@@
By decomposing the particle velocity into a mean flow with fluctuations, we find that the macroscopic velocity $\mathbf{u}$ is governed by the mean phase gradient $\nabla\Psi$ scaled by the local order parameter $R$: $\mathbf{u}\approx R\nabla\Psi$. While we derive a set of constraining equations that are isomorphic to shallow water hydrodynamics, there are striking differences: the effective inertial mass in our model scales with synchronization:
$\lambda=R^{2}.$ (6)
At the same time, the effective pressure scales with disorder:
$P_{\text{eff}}=\frac{v_{0}^{2}}{2}(1-R^{4}).$ (7)
Crucially, a stiff active pressure enforces a sound speed proportional to order ($c_{s}\propto R$). Since the inertial velocity also scales with $R$, this locks the flow into an inherently supersonic state (Mach number >1>1). Consequently, defect cores ($R\rightarrow 0$) are isolated behind Mach cones, creating sonic horizons that prevent radiative decay. While fast compressible shocks disperse inertial energy locally, the time-averaged background flow remains incompressible. This effectively traps energy within the topological degrees of freedom, meaning the macroscopic dynamics are governed strictly by the conservation of topological charge.
@@ VII.2 Momentum and Pressure
confirming that the effective inertial mass $\lambda$ scales with the square of the order parameter. Subsequently, matching the isotropic (constant) terms determines the effective pressure:
$\frac{v_{0}^{2}}{2}=P_{\text{eff}}+\lambda\frac{v_{0}^{2}R^{2}}{2}.$ (30)
Substituting $\lambda=R^{2}$ yields,
$P_{\text{eff}}=\frac{v_{0}^{2}}{2}-\frac{v_{0}^{2}R^{4}}{2}=\frac{v_{0}^{2}}{2}(1-R^{4}).$ (31)
Thus, we recover an effective pressure, generated by the incoherent, random fluctuations of the particles that are not aligned with the macroscopic flow, that scales with $\propto(1-R^{4})$.
It is pertinent to compare this result to the swim pressure derived in the kinetic theory of active fluids (e.g., [67]), which typically scales with the velocity variance as $\propto(1-R^{2})$. The discrepancy arises from the macroscopic transport. In standard active fluids, the particle mass is constant (corresponding to $\lambda=1$), meaning the coherent kinetic energy scales as $u^{2}\propto R^{2}$. In our topological gas, the effective inertial mass itself depends on local synchronisation ($\lambda=R^{2}$), causing the “coherent energy” to scale as $\lambda u^{2}\propto R^{4}$.
Consequently, our effective pressure $P_{\text{eff}}\propto(1-R^{4})$ decays much more slowly with ordering than the standard $(1-R^{2})$ prediction; implying a much stiffer equation of state near ordered regions ($R\approx 1$) than standard active fluids. In our model, this creates a thermodynamic barrier: the isotropic ’disorder pressure’ remains dominant even as the system attempts to synchronize. This explains the stability of the active vortex glass (Phase II): the intermediate agitation ($\Delta\omega$) is insufficient to overcome this pressure plateau, causing the system to jam in a disordered state ($R\to 0$). Only the high agitation of Phase III provides the energy density required to penetrate this barrier and sustain the regime of marginal synchronization ($0<R<1$) characteristic of the Onsager condensate.
Returning to the task at hand, we substitute the closure back into the momentum conservation equation and obtain,
@@
VII.3 The shallow water analogy
As stated in the Methodology section, the interpretation of Eq. (18) identifies the system being modeled as isomorphic to the shallow water framework, where the active density $\rho$ acts as the fluid height.
However, before a rigorous derivation and scrutiny of the analogy, we must point out a nuance introduced by a critical thermodynamic distinction in the equation of state. In classical shallow water systems, hydrostatic pressure scales with height ($P\propto\rho^{2}$), leading to dispersion [65]. In Eq. (31), we derived the pressure term $P_{\text{eff}}\propto(1-R^{4})$, meaning that the effective pressure scales with disorder. Crucially, because the local order parameter $R$ arises from the mean-field convolution of the particle density (Eq. 1), regions of high density support high synchronisation ($R\sim\rho$; realized by $R$ being the magnitude of the complex mean field, in turn a measurement of $\rho$, since phasors align in local synchronisation regions). Consequently, the pressure gradient $\nabla P_{\text{eff}}\propto-\nabla R^{4}$ acts as a suction force, drawing agents from disordered regions ($R\approx 0$) toward ordered vortex cores ($R\approx 1$). This manifests as a fluid with effective negative compressibility, driving the system to phase-separate into the coherent topological structures observed in the Onsager dipole, rather than dispersing into noise.
Consequently, the inviscid Euler behavior that we observe in the RFE field represents the low-Mach-number limit of this compressible system, where the rotational, non-compressive wave modes dominate the irrotational, compressive wave (acoustic) modes at large scales. Hence, in the hydrodynamic limit (the Onsager dipole, or Phase III in Table 1), we recover in Eq. (32) the effective Euler advection term $\lambda(\mathbf{u}\cdot\nabla)\mathbf{u}$. Here we note that the vanishing inertia ($\lambda\to 0$) at defect cores allow for instantaneous acceleration of defects in response to potential gradients, which mimics an infinite sound speed ($c_{s}\to\infty$), physically justifying the rigid lid assumption by allowing the flow field to adjust without acoustic delay.
It is instructive at this point to contrast the foregoing with the defect-driven accumulation observed in active nematics [68] and flow-stabilized vortex lattices [69]. In those systems, mass transport is driven by the divergence of an active stress that scales linearly with the nematic tensor order parameter ($\sigma\sim\mathbf{Q}$), which creates effective pressure wells that scale akin to $\propto(1-R^{2})$. In our model, however, the emergent inertia of the topological gas ($\lambda=R^{2}$) yields a stiffer effective pressure $P_{\text{eff}}\propto(1-R^{4})$ (Eq. 31). This creates sharper, ’hard-core’ density cavities at defect sites (see Figure 7 below), providing additional justification for the point vortex (Kirchhoff-Onsager) Hamiltonian in topological gas dynamics, since it assumes distinct, non-overlapping points [22].
VII.4 The Active Enthalpy Functional
@@
In this section, we shall rigorously close the derivation of the shallow water analogy, using the active enthalpy functional as a theoretical vehichle [70, 71]. However, we must first bolster the connection with a specific empirical observation. Figure 7 shows the end-state of a Onsager dipole simulation, color-coding $\omega_{i}$ (top), $R$ (middle), and topological charge $m$ (bottom). We make two important observations;
@@
where $g_{\text{eff}}$ is the effective coupling constant representing the stiffness of the order parameter (analogous to gravity $g$ in shallow waters), and $\gamma$ is a coupling coefficient with dimensions $\text{length}^{2}\text{time}^{-1}$ relating frequency disorder $\omega(\mathbf{r})$ to potential energy. Two concise observations can now be had by dimensional analysis. (1) the units of $\gamma$ naturally lead to the definition $\gamma=\sigma v_{0}$, anchoring the topographic potential to the physical interaction scale of the agents. (2) by direct comparison between Eqs. (31) and (33), we observe that $g_{\text{eff}}\equiv v_{0}^{2}/2$.
@@
VII.5 Active acoustic speed
At this point, we must address acoustics, as a shallow water vortex diffuses into the medium through sound waves. To demonstrate why our Onsager dipole is unable to do so, we first revisit the continuity and momentum equations (Eqs. 17 and 37) using $\rho=R^{2}$,
@@ VII.5 Active acoustic speed
VII.6 Hydrodynamic closure
Finally, we reconcile the stiffness of the effective pressure ($P_{\text{eff}}\propto 1-R^{4}$) with the observation of compressible shocks. As the active acoustic speed (Eq. 46) scales linearly with order ($c_{s}\propto R$), the stiffness ensures that the sound speed inherently lags behind the inertial flow velocity ($|\mathbf{u}|\propto R$, Eq. 14), forcing the system into a global supersonic state with $M\approx\sqrt{2}$. As we discussed in the foregoing subsection, this condition mandates the formation of synchronization shocks (Mach cones) at defect sites, which act as the primary mechanism topological sorting (yielding Eulerian Onsager-Kirchoff dynamics in the supersonic topological gas). This leads us to assert the hazardous rigid lid incompressibility,

### v5

@@ II Methodology
At this point, We identify the characteristic scales in out model; the kernel size $\sigma$ and the active swim speed $v_{0}$, which yield the interaction time $\tau\equiv\sigma/v_{0}=(3\text{ grid units})/(0.5\text{ grid units/s})=6$ s, yielding the dimensionless units $\tilde{x}=x/\sigma$, $\tilde{t}=t/\tau$, and $\tilde{\omega}=\omega(\sigma/v_{0})$ (we subsequently drop the tilde). The system size becomes $\Lambda=2L/\sigma\approx 13$. A natural definition of the RFE is then afforded by $\tau$; we define the RFE as the center of mass of a particle’s trajectory during $\tau$, effectively course-graining the singularities:
@@
By decomposing the particle velocity into a mean flow with fluctuations, we find that the macroscopic velocity $\mathbf{u}$ is governed by the mean phase gradient $\nabla\Psi$ scaled by the local order parameter $R$: $\mathbf{u}\approx R\nabla\Psi$. While we derive a set of constraining equations that are isomorphic to shallow water hydrodynamics, there are striking differences: the effective inertial mass in our model scales with order:
$\lambda=R^{2},$ (6)
with an effective restoring potential (enthalpy),
$h_{\text{eff}}=\frac{v_{0}^{2}}{2}R^{2}.$ (7)
The macroscopic restoring force is governed by the Eq. (7)’s gradient, enforcing a sound speed that is also proportional to order,
$c_{s}=\frac{v_{0}R}{\sqrt{2}}.$ (8)
Since the inertial velocity likewise scales with $R$, this locks the flow into an inherently supersonic state (Mach number >1>1). Consequently, defect cores ($R\rightarrow 0$) are isolated behind Mach cones, creating sonic horizons that prevent radiative decay. While fast compressible shocks disperse inertial energy locally, the time-averaged background flow remains incompressible. This effectively traps energy within the topological degrees of freedom, meaning the macroscopic dynamics are governed strictly by the conservation of topological charge.
@@ VII.2 Momentum and Pressure
confirming that the effective inertial mass $\lambda$ scales with the square of the order parameter.
It is pertinent to compare this result to the swim pressure derived in the kinetic theory of active fluids (e.g., [67]). In standard active fluids, the particle mass is constant (corresponding to $\lambda=1$), meaning the coherent kinetic energy scales as $u^{2}\propto R^{2}$. In our topological gas, the effective inertial mass itself depends on local synchronisation ($\lambda=R^{2}$), causing the “coherent energy” to scale as $\lambda u^{2}\propto R^{4}$.
Next, we substitute the closure back into the momentum conservation equation and obtain,
@@
VII.3 Shallow waters and the active enthalpy functional
@@
As stated in the Methodology section, the interpretation of Eq. (19) identifies the system being modeled as isomorphic to the shallow water framework, where the active density $\rho$ acts as the fluid height. To proceed, we shall use the active enthalpy functional as a theoretical vehichle [68, 69]. However, we must first bolster the connection with a specific empirical observation. Figure 7 shows the end-state of a Onsager dipole simulation, color-coding $\omega_{i}$ (top), $R$ (middle), and topological charge $m$ (bottom). We make two important observations;
@@
where $g_{\text{eff}}$ is the effective coupling constant representing the stiffness of the order parameter (analogous to gravity $g$ in shallow waters), and $\gamma$ is a coupling coefficient with dimensions $\text{length}^{2}\text{time}^{-1}$ relating frequency disorder $\omega(\mathbf{r})$ to potential energy. Two concise observations can now be had by dimensional analysis. (1) the units of $\gamma$ naturally lead to the definition $\gamma=\sigma v_{0}$, anchoring the topographic potential to the physical interaction scale of the agents. (2) by direct comparison between Eqs. (LABEL:eq:peff2) and (32), we observe that $g_{\text{eff}}\equiv v_{0}^{2}/2$.
@@
VII.4 Active acoustic speed
At this point, we must address acoustics, as a shallow water vortex diffuses into the medium through sound waves. To demonstrate why our Onsager dipole is unable to do so, we first revisit the continuity and momentum equations (Eqs. 18 and 36) using $\rho=R^{2}$,
@@ VII.4 Active acoustic speed
VII.5 Hydrodynamic closure
As the active acoustic speed (Eq. 45) scales linearly with order ($c_{s}\propto R$), the sound speed inherently lags behind the inertial flow velocity by a constant, order unity, forcing the system into a global supersonic state with $M\approx\sqrt{2}$. As we discussed in the foregoing subsection, this condition mandates the formation of synchronization shocks (Mach cones) at defect sites, which act as the primary mechanism topological sorting (yielding Eulerian Onsager-Kirchoff dynamics in the supersonic topological gas). This leads us to assert the hazardous rigid lid incompressibility,